OpenStudy (anonymous):

Definition, Property, Postulate, or Theorem that justifies. If XY=WZ the XY+TU=WZ+TU I need the name of the theorem or postulate.

Nnesha (nnesha):

hint :they adding TU both side

OpenStudy (mathstudent55):

When you solve an equation, such as this one: x - 2 = 5 by doing this: x - 2 + 2 = 5 + 2 (adding 2 to both sides) What allows you to add the same number to both sides?

OpenStudy (anonymous):

so what theorem?

OpenStudy (mathstudent55):

You are adding the same amount to both sides. It is the Addition Property of Equality.

Nnesha (nnesha):

you are adding so it relates to add

OpenStudy (anonymous):

thx

OpenStudy (mathstudent55):

yw